OpenStudy (anonymous):

polynomial function that has zeros at -3,2, and 4

OpenStudy (campbell_st):

if x = -3 is a zero, then x + 3 is a factor x =2 is a zero, then x -2 is a factor x = 4 is a zero, then x - 4 is a factor so that means f(x) =(x +3)(x-2)(x-4) just distribute